# Spin noise spectroscopy of an alignment-based atomic magnetometer

### v1

@@ II.3 Experimental setup
Figure 2 shows a schematic of the experimental setup we employ. Linearly polarised light, with a wavelength of 895 nm, is passed through an optical fibre. The light is resonant with the $F=4\rightarrow F^{\prime}=3$ D1 transition of caesium. The light has an electric field amplitude $\bm{E}_{0}$ and passes through a cubic paraffin-coated cell containing caesium atoms, which is placed inside a magnetic shield (Twinleaf MS-1). The cell is $(5\text{mm})^{3}$ and is kept at room temperature ($\sim 18^{\circ}$).
@@
A static magnetic field $\bm{B}_{0}=B_{0}\,\bm{e}_{z}$, with $B_{0}=2.74~{}\mu$T, is applied using the magnetic shield coils. This corresponds to inducing Larmor precession of the atomic spin at a frequency of $f_{L}\approx 10$ kHz. A half-wave plate is placed before the magnetic shield to change the angle of the electric field amplitude of the light, $\bm{E}_{0}$, with respect to the direction of the static magnetic field. Consistently with Fig. 1(c), we denote in Fig. 2 by $\theta$ the angle between the linear polarisation and the $z$-direction in the $yz$-plane, so that, e.g., $\bm{E}_{0}=E_{0}\,\bm{e}_{z}$ for $\theta=0$ and $\bm{E}_{0}=E_{0}\,\bm{e}_{y}$ for $\theta=\pi/2$. In parallel, a noisy magnetic field $\bm{B}_{\mathrm{noise}}(t)=B_{\mathrm{noise}}(t)\,\bm{e}_{x}$ with $B_{\mathrm{noise}}(t)$ being varied in the experiment from $0.132~{}\mathrm{pT_{rms}}$ to $32.92~{}\mathrm{nT_{rms}}$ is applied to the system using a function generator connected to a homemade square Helmholtz coil.
@@
As shown in Fig. 2, after the magnetic shield the polarisation rotation of the transmitted light is measured using a half-wave plate ($\lambda/2$) and a polarising beam splitter (PBS). The half-wave plate is rotated to match exactly the polarisation angle $\theta^{\prime}$ of the transmitted light, see Fig. 1(c), so that it is the deviations from $\theta^{\prime}$ that are then effectively measured via balanced photodection (BPD)  [^2] Even if $\theta^{\prime}$ is not exactly matched, the resulting DC-component of the detected signal (27) yields a spike at zero frequency in the PSD, whose presence may be safely ignored within our analysis.. In particular, the difference in the intensity of the outgoing beams from the PBS is tracked using a Thorlabs balanced photodetector (PDB210A/M). The output photocurrent is recorded in real time using a data acquisition card. Further technical details on the magnetometer can also be found in Ref. [25].
Figure 2: Alignment-based magnetometer: experimental setup. The laser light propagates along the $x$-direction, while the angle $\theta$ of its linear polarisation can be adjusted by rotating half-wave plates ($\lambda/2$) before it enters the magnetic shield (marked in grey). Within the shield, the light simultaneously pumps and probes the atomic ensemble of caesium atoms contained within a paraffin-coated cell (Cell). The atoms are subject to (strong) static and (relatively weak) noisy magnetic fields: $\bm{B}_{0}$ applied in the $z$-, and $\bm{B}_{\mathrm{noise}}$ applied in the $x$-direction, respectively. Due to the Faraday effect, the polarisation of the light exhibits rotation, which is then measured by passing the output beam through a polarising beam splitter (PBS) and performing balanced photodetection (BPD).
@@
Importantly, as our experiment matches the spatial configuration of an alignment-based magnetometer described in previous sections, we can interpret the measured photocurrent of the BPD exactly as $\delta S(t)\coloneqq S(t)-\left\langle S(t)\right\rangle$, i.e. as deviations from the mean value of the detected signal stated in Eq. (27). The effective proportionality constant $g_{D}$ in Eq. (27), which relates the instantaneous atomic state to the photocurrent signal, is then dictated by a number of experimental conditions including: the light power (1 $\mu W$), pumping efficiency, size of the vapour cell ((5 mm)${}^{3}$), as well as the temperature (room temperature, $\sim 18^{\circ}$). On the other hand, the detection noise, $\zeta(t)$ in Eq. (27), should be attributed to photon shot-noise and electronic jitter arising solely due to the photo-detection process that effectively leads to a background noise—within the measured PSD a DC offset is observed independently whether the light beam interacts with the atoms or not. This results in a noise floor that depends on the frequency, partially due to the 1/f noise. This will be taken into account when interpreting the data in Sec. III.2 below.
@@ III Spin noise spectroscopy
Denoting the Fourier transform of any signal, here the measured current of the balanced photodetector $\delta S(t)$, over a finite-time interval $[0,T]$ as:
@@
[…]
where by “ss”, as before, we emphasise the above to hold in the steady state. Crucially, Eq. (30) is now expressed in terms of the auto-covariance function of the detected signal, $S(t)$ specified in Eq. (27), i.e. $\left\langle S(t_{1}),S(t_{2})\right\rangle\coloneqq\left\langle S(t_{1})S(t_{2})\right\rangle-\left\langle S(t_{1})\right\rangle\left\langle S(t_{2})\right\rangle$ 333 Eq. (30) could be equivalently defined in terms of the auto-correlation function of $\delta S(t)=S(t)-\left\langle S(t)\right\rangle$ [56]. .
Substituting explicitly the form of the detected signal (27) into Eq. (30), we obtain the form of the PSD applicable to our problem as
@@
with $p,q=-2,..,2$ specifying the coefficients of the $\textbf{m}_{t}$-vector (20) evaluated in the steady state. As expected, the detection noise $\zeta(\omega)$, being uncorrelated from any other noise sources, leads to just a noise floor in Eq. (29).
@@ III.1 Theoretical predictions
$\gamma_{2}$ $\approx\Gamma_{2}+\frac{1}{2}\omega_{\mathrm{noise}}.$ (37)
All $\omega_{j}$ and $\gamma_{|j|}$ stated in Eqs. (36) and (37), respectively, can be determined analytically, however, we already simplified their forms above for $\omega_{\pm 2}$, $\gamma_{0}$ and $\gamma_{2}$, which are given $\forall_{j}:\;\Omega_{L}\gg\Gamma_{|j|}$ and $\Omega_{L}\gg\omega_{\mathrm{noise}}$.
@@
The dispersive equivalents of expressions (III.1) can also be determined analytically (given $\forall_{j}:\;\Omega_{L}\gg\Gamma_{|j|}$ and $\Omega_{L}\gg\omega_{\mathrm{noise}}$) and can be found in App. C. However, see App. C.1, these are irrelevant upon substituting the parameters applicable to our experiment. Hence, we ignore their contribution to the PSD (35) from now on.
Although Eqs. (III.1) allow us to predict the dependence of the peak heights for all values of noise intensity $\omega_{\mathrm{noise}}$, it directly follows that for low noise-strengths they obey $p_{0}^{a}~{}\propto~{}\omega_{\mathrm{noise}}^{2}$, $p_{1}^{a}\propto\omega_{\mathrm{noise}}$, and $p_{2}^{a}\propto\omega_{\mathrm{noise}}^{2}$. Furthermore, their angular dependence factorises and is given by
@@
[…]
Figure 4: Amplitudes (a) and line-widths (b) of the PSD peaks as a function of the white-noise strength, which is varied from 7 $\mathrm{mV_{rms}}$ to 3500 $\mathrm{mV_{rms}}$ with the light polarisation angle fixed to $\theta=25^{\circ}$. For every white-noise value $V_{\mathrm{noise}}$, amplitudes and line-widths are determined for each of the three relevant (absorptive) peaks in the PSD via the fitting procedure described in Fig. 3. Using the same colouring as in the main plot of Fig. 3, the theoretical predictions are shown for each peak at 00 (red, dot-dashed), $\approx f_{L}$ (green, solid) and $\approx 2f_{L}$ (blue, dashed) frequency. (a) All three peak amplitudes $\widetilde{p}_{j}^{a}$ clearly follow the theoretical predictions of Eq. (III.1), which we fit by tuning only the overall proportionality constants ($C_{j}$ for each $j$, see main text) while using the values of dissipation rates $\widetilde{\bm{\Gamma}}$ determined from the line-width data. (b) As predicted by Eq. (37), the line-widths increase linearly with $f_{\mathrm{noise}}$($=cV_{\mathrm{noise}}^{2}$) with proportionality constants agreeing almost exactly for each peak. The line-widths at $V_{\mathrm{noise}}=0$ provide then the corresponding dissipation rates $\widetilde{\bm{\Gamma}}$ (in Hz): $\widetilde{\Gamma}_{0}=27(3)$, $\widetilde{\Gamma}_{1}=33(3)$ and $\widetilde{\Gamma}_{2}=30(3)$.
@@ III.2 Measured noise spectra
In order to validate the theoretical model outlined above, we vary the noisy magnetic field in the table-top alignment-based magnetometer described in Sec. II.3. For the purpose of experiment, we define the noise spectral density in Hz, i.e. $f_{\mathrm{noise}}$ such that $\omega_{\mathrm{noise}}=2\pi f_{\mathrm{noise}}$. As we apply a known voltage through the coil, it is convenient to further write $f_{\mathrm{noise}}=c\,V_{\mathrm{noise}}^{2}$, where $V_{\mathrm{noise}}^{2}$ has now units of $\mathrm{mV_{rms}}$ and the proportionality constant $c$ can be explicitly determined for our setup, see App. D.
[…]
For such a low strength of white noise the three peaks are clearly separated and effectively experience detection noise (and $1/f$-noise) of different strengths. Hence, we fit the absorptive Lorentzian functions appearing in Eq. (35) to each of them individually, i.e. we fit independently around each peak with $j=0,1,2$ the profile:
$\mathrm{PSD}_{j}(f)=\frac{\widetilde{p}_{j}^{a}\widetilde{\gamma}_{j}^{2}}{(f-f_{j})^{2}+\widetilde{\gamma}_{j}^{2}}+\left\langle|\widetilde{\zeta}_{j}(f)|^{2}\right\rangle,$ (42)
where the central frequencies $f_{j}\approx jf_{L}$ (with $f_{0}=0$ imposed) and the line widths $\widetilde{\gamma}_{j}=\gamma_{j}/2\pi$ are now expressed in Hz, while both the peak amplitude and noise floor, $\widetilde{p}_{j}$ and $\left\langle|\widetilde{\zeta}_{j}(f)|^{2}\right\rangle$ respectively, have units of $\mathrm{V}^{2}/\mathrm{Hz}$.
$j$
$f_{j}$
$\widetilde{\gamma}_{j}$
$\widetilde{p}_{j}^{a}$
$\left\langle|\widetilde{\zeta}(f)|^{2}\right\rangle$
[…]
00
00
$34.3(3)$
$9.15(5)\times 10^{-12}$
$6.5(8)\times 10^{-14}$
$1$
$9603.1(3)$
$48.9(4)$
$4.17(2)\times 10^{-13}$
$6.35(4)\times 10^{-14}$
$2$
$19208.0(3)$
$37.9(4)$
$1.97(1)\times 10^{-13}$
$2.67(2)\times 10^{-14}$
[…]
In contrast, when dealing with high strengths of white noise ($V_{\mathrm{noise}}\geq 1~{}\mathrm{V_{rms}}$), the three peaks strongly overlap within the PSD, as shown explicitly for $V_{\mathrm{noise}}=3~{}\mathrm{V_{rms}}$ within the inset of Fig. 3. As predicted by the theory, this is due to an apparent increase of the peak line-widths. Moreover, as the peak amplitudes are then at least an order of magnitude above the noise floor, the latter may be assumed to be frequency-independent. As a consequence, it is then correct to fit a single function to the whole PSD, i.e. the complete absorptive part of Eq. (35).
@@
In what follows, we verify in more detail the expressions (III.1) for the peak amplitudes by studying explicitly their dependence on the noise spectral density, $f_{\mathrm{noise}}$, and the light polarisation angle, $\theta$.

### v2

@@ II.3 Experimental setup
Figure 2 shows a schematic of the experimental setup we employ. Linearly polarised light, with a wavelength of 895 nm, is passed through a polarization-maintaining optical fiber. The light is resonant with the $F=4\rightarrow F^{\prime}=3$ D1 transition of caesium. The light has an electric field vector $\bm{E}_{0}$ and passes through a cubic paraffin-coated cell containing caesium atoms, which is placed inside a magnetic shield (Twinleaf MS-1). The cell is $(5\text{mm})^{3}$ and is kept at room temperature ($\sim 18^{\circ}$). The laser beam diameter is $\sim 2$ mm.
@@
A static magnetic field $\bm{B}_{0}=B_{0}\,\bm{e}_{z}$ is applied using the magnetic shield coils. A half-wave plate is placed before the magnetic shield to change the angle of the electric-field vector of the light, $\bm{E}_{0}$, with respect to the direction of the static magnetic field. Consistently with Fig. 1(c), we denote in Fig. 2 by $\theta$ the angle between the linear polarisation and the $z$-direction in the $yz$-plane, so that, e.g., $\bm{E}_{0}=E_{0}\,\bm{e}_{z}$ for $\theta=0$ and $\bm{E}_{0}=E_{0}\,\bm{e}_{y}$ for $\theta=\pi/2$. In parallel, a noisy magnetic field $\bm{B}_{\mathrm{noise}}(t)=B_{\mathrm{noise}}(t)\,\bm{e}_{x}$ is applied to the system using a home-made square Helmholtz coil. The current driving the coil corresponds to the white-noise signal outputted by a function generator, which is, however, first transformed thorough a $f_{\textrm{cutoff}}=1$ MHz low-pass filter to prevent the effect of aliasing, as well as a $\delta\!f=1$ kHz high-pass filter to eliminate any spurious contributions at very low frequencies (see App. D for more details).
When measuring the power spectra of our magnetometer as a function of the strength of the noisy magnetic field, as described below, we set the static magnetic field to $B_{0}=2.7~{}\mu$T, what corresponds to inducing the Larmor precession of the atomic spin at a frequency $f_{L}\approx 9.45$ kHz. In such a case, each dataset is acquired for a given fixed white-noise amplitude in the range of $70.91~{}\mathrm{nT_{rms}}$, for the first set, to $35.45~{}\mathrm{\mu T_{rms}}$, for the final dataset. When investigating the dependence on the polarisation angle $\theta$ instead, we use the static field $B_{0}=2.74~{}\mu$T (i.e. $f_{L}\approx 9.6$ kHz), while the white-noise amplitude is fixed to $1.4~{}\mathrm{\mu T_{rms}}$.
@@
As shown in Fig. 2, after the magnetic shield the polarisation rotation of the transmitted light is measured using a half-wave plate ($\lambda/2$) and a polarising beam splitter (PBS). The linear polarisation of the set up was checked before and after the vapor cell, by measuring the ellipticity of the beam which was found to be of negligible order. The half-wave plate is rotated to match exactly the polarisation angle $\theta^{\prime}$ of the transmitted light, see Fig. 1(c), so that it is the deviations from $\theta^{\prime}$ that are then effectively measured via balanced photodection (BPD) [^2] Even if $\theta^{\prime}$ is not exactly matched, the resulting DC-component of the detected signal (27) yields a spike at zero frequency in the PSD, whose presence may be safely ignored within our analysis.. In particular, the difference in the intensity of the outgoing beams from the PBS is tracked using a Thorlabs balanced photodetector (PDB210A/M). The output photocurrent is recorded in real time using a data acquisition card. Further technical details about the performance of the magnetometer used can be found in Ref. [25].
Importantly, as our experiment matches the spatial configuration of an alignment-based magnetometer described in previous sections, we can interpret the measured photocurrent of the BPD exactly as:
$\delta S(t)\coloneqq S(t)-\bar{S},$ (28)
@@
where by $\bar{S}\coloneqq\left\langle S(t)\right\rangle_{\mathrm{ss}}$ we denote the time-independent mean DC component of the measured signal, which is determined by the mean of the steady-state solution of Eq. (24). As a result, Eq. (28) describes deviations from the mean value of the detected signal stated in Eq. (27). The effective proportionality constant $g_{D}$ in Eq. (27), which relates the instantaneous atomic state to the photocurrent signal, is then dictated by a number of experimental conditions including: the light power (1 $\mu W$), pumping efficiency, size of the vapour cell ((5 mm)3), as well as the temperature (room temperature, $\sim 18^{\circ}$). On the other hand, the detection noise, $\zeta(t)$ in Eq. (27), should be attributed to photon shot-noise and electronic jitter arising solely due to the photo-detection process that effectively leads to a background noise—within the measured PSD a DC offset is observed independently whether the light beam interacts with the atoms or not. This results in a noise floor that depends on the frequency, partially due to the 1/f noise. This will be taken into account when interpreting the data in Sec. III.2 below.
@@ III Spin noise spectroscopy
Denoting the Fourier transform of any signal, here the measured current of the balanced photodetector $\delta S(t)$ defined in Eq. (28), over a finite-time interval $[0,T]$ as:
@@
$\mathrm{PSD}(\omega)\;$ =ss $\underset{\textrm{ss}}{=}\;2\int_{0}^{\infty}\mathrm{d}t\cos(\omega t)\left\langle\delta S(t)\delta S(0)\right\rangleunder$ (31a)
=ss $\underset{\textrm{ss}}{=}\;2\int_{0}^{\infty}\mathrm{d}t\cos(\omega t)\left\langle S(t),S(0)\right\rangleunder$ (31b)
where by “ss”, as before, we emphasise the above to hold in the steady state. In particular, see Eq. (31a), the PSD (30) can be expressed in terms of the auto-correlation function of the (zero-mean) signal $\delta S(t)$; or equivalently, see Eq. (31b), the auto-covariance function of the actual detected signal, $S(t)$ specified in Eq. (27), i.e. $\left\langle S(t_{1}),S(t_{2})\right\rangle\coloneqq\left\langle S(t_{1})S(t_{2})\right\rangle-\left\langle S(t_{1})\right\rangle\left\langle S(t_{2})\right\rangle$ [59].
Substituting explicitly the form of the detected signal (27) into Eq. (31b), we obtain the form of the PSD applicable to our problem as
@@
with $p,q=-2,..,2$ specifying the coefficients of the $\textbf{m}_{t}$-vector (20) evaluated in the steady state. As the detection noise $\zeta(\omega)$ is uncorrelated from any other noise sources within our model, it leads to a noise floor in Eq. (30) as expected.
@@ III.1 Theoretical predictions
$\gamma_{2}$ $\approx\Gamma_{2}+\frac{1}{2}\omega_{\mathrm{noise}},$ (38)
with $\Gamma_{i}$ defined as in Eq. (23). All $\omega_{j}$ and $\gamma_{|j|}$ stated in Eqs. (37) and (38), respectively, can be determined analytically, however, we already simplified their forms above for $\omega_{\pm 2}$, $\gamma_{0}$ and $\gamma_{2}$, which are given by $\forall_{j}:\;\Omega_{L}\gg\Gamma_{|j|}$ and $\Omega_{L}\gg\omega_{\mathrm{noise}}$.
@@
The dispersive equivalents of expressions (39) can also be determined analytically (given $\forall_{j}:\;\Omega_{L}\gg\Gamma_{|j|}$ and $\Omega_{L}\gg\omega_{\mathrm{noise}}$) and can be found in App. C. However, see App. C.1, these are negligible upon substituting the parameters applicable to our experiment. Hence, we ignore their contribution to the PSD (36) from now on.
Although Eqs. (39) allow us to predict the dependence of the peak heights for all values of noise intensity $\omega_{\mathrm{noise}}$, it directly follows that for low noise-strengths they obey $p_{0}^{a}~{}\propto~{}\omega_{\mathrm{noise}}^{2}$, $p_{1}^{a}\propto\omega_{\mathrm{noise}}$, and $p_{2}^{a}\propto\omega_{\mathrm{noise}}^{2}$. Furthermore, their angular dependence factorises and is given by
@@
Consistently, these angular dependencies correspond to the squares of the expressions derived in Ref. [61], when the response to a radio-frequency magnetic field is considered instead of white noise.
@@ III.2 Measured noise spectra
In order to validate the theoretical model outlined above, we vary the noisy magnetic field in the table-top alignment-based magnetometer described in Sec. II.3. For the purpose of experiment, we define the noise spectral density in Hz, i.e. $f_{\mathrm{noise}}$ such that $\omega_{\mathrm{noise}}=2\pi f_{\mathrm{noise}}$. As we apply a known voltage through the coil, it is convenient to further write $f_{\mathrm{noise}}=c\,V_{\mathrm{noise}}^{2}$, where $V_{\mathrm{noise}}$ has units of $\mathrm{mV_{rms}}$ and the proportionality constant $c$ can be explicitly determined for our setup, see App. D.
Figure 3: Power spectral density (PSD) measured in the experiment against the theoretical fits predicted by Eq. (36). The signal was recorded over 100 time traces, each of 1 second duration. Within the main plot, relatively low-noise spectral density of $f_{\mathrm{noise}}=0.26$ Hz is chosen, corresponding to 140 mVrms (the input polarisation angle is set to $\theta=40^{\circ}$). The so-obtained PSD has the experimental noise floor subtracted and is then fitted using a single function (magenta curve) containing three Lorentzian peaks with centres located close to the frequencies: $f=0$, $f_{L}$ and $2f_{L}$. Here, the three peaks are clearly distinguishable due to their small linewidths. In contrast, the inset shows the PSD for high-noise spectral density $f_{\mathrm{noise}}=120$ Hz, corresponding to 3 Vrms (with the input polarisation angle set now to $\theta=25^{\circ}$), that nonetheless is well described by the same fit-function (yellow curve) despite the Lorentzian peaks now overlapping significantly.
In order to measure the PSD (30) of our device, we record for a given value of $f_{\mathrm{noise}}$ one hundred one-second datasets, for each of which the Fourier transform is then computed before averaging. The main plot of Fig. 3 shows an exemplary PSD obtained for the input polarisation angle being set to $\theta=40^{\circ}$, and a relatively low noise spectral density being applied, $f_{\mathrm{noise}}=0.26$ Hz (corresponding to $V_{\mathrm{noise}}=140$ mVrms). It is clear that the PSD has peaks at approximately 00, $f_{L}$ and $2f_{L}$ frequencies with $\Omega_{L}=2\pi f_{L}$, as anticipated by Eq. (37).
Due to a clear separation of the peaks, the noise floor varies between them—it takes the value of approximately $6.5\times 10^{-14}$ V2/Hz for the peaks centred at $f_{0}=0$ and $f_{1}\approx f_{L}$, while it reads about $2.7\times 10^{-14}$ V2/Hz for the peak at $f_{2}\approx 2f_{L}$. When fitting the data we subtract the noise floor (data not shown) of the experiment from each data set. This removes the detection noise (and $1/f$-noise) of different strengths across the spectrum. We then fit a single function containing three Lorentzian peaks to the whole spectrum, i.e. the complete absorptive part of Eq. (36). The corresponding fit parameters obtained for each of the three peaks are listed in Table 1. In contrast, when dealing with high strengths of white noise, the three peaks strongly overlap within the PSD, as shown explicitly for $f_{\mathrm{noise}}=120~{}\mathrm{Hz}$ within the inset of Fig. 3. As predicted by the theory, this is due to an apparent increase of the peak linewidths.
Figure 4: Amplitudes (a) and linewidths (b) of the PSD peaks as a function of the white-noise strength, which is varied from $6.53\times 10^{-4}$ Hz to 163 Hz with the light polarisation angle fixed to $\theta=25^{\circ}$. For every white-noise value $f_{\mathrm{noise}}$, amplitudes and linewidths are determined for each of the three relevant (absorptive) peaks in the PSD via the fitting procedure described in Fig. 3. The theoretical predictions are shown for each peak at 00 (red, dot-dashed), $\approx f_{L}$ (green, solid) and $\approx 2f_{L}$ (blue, dashed) frequency. (a) All three peak amplitudes $\widetilde{p}_{j}^{a}$ clearly follow the theoretical predictions of Eq. (39), which we fit by tuning the overall proportionality constants (allowing different $C_{j}$ for each $p^{a}_{j}$ in Eq. (39)) and common values of the dissipation rates ($\Gamma_{0}$, $\Gamma_{1}$ and $\Gamma_{2}$ in Eq. (39)). (b) The linewidths increase linearly with $f_{\mathrm{noise}}$ with the slope agreeing almost exactly for each peak with the proportionality constants predicted by Eq. (38). Moreover, their offsets at $f_{\mathrm{noise}}=0$ provide us independently with the dissipation rates ($\Gamma_{0}$, $\Gamma_{1}$ and $\Gamma_{2}$ appearing also in Eq. (38)), which are consistent (up to $\approx\!10$Hz) with their equivalents predicted by amplitude-fitting in (a)—see the main text.
@@
In what follows, we verify in more detail the expressions (39) for the peak amplitudes by studying explicitly their dependence on the noise spectral density, $f_{\mathrm{noise}}$, and the light polarisation angle, $\theta$.
$j$ $f_{j}$ $\widetilde{\gamma}_{j}$ $\widetilde{p}_{j}^{a}$
(Hz) (Hz) (V2/Hz)
00 00 $34.9(2)$ $9.03(3)\times 10^{-12}$
$1$ $9603(1)$ $48.6(4)$ $4.16(2)\times 10^{-13}$
$2$ $19208(1)$ $37.8(6)$ $1.97(2)\times 10^{-13}$
Table 1: Fit parameters in the low-noise regime. A single function corresponding to the absorptive part of Eq. (36) is fitted to reconstruct the PSD presented in the main plot of Fig. 3. As it contains three Lorentzian profiles ($j=0,1,2$), such a procedure yields three distinct sets of parameters: central frequencies ($f_{j}$), linewidths ($\widetilde{\gamma}_{j}$) and peak amplitudes ($\widetilde{p}_{j}^{a}$). The data was collected in the above measurement of the PSD (main plot of Fig. 3) for Larmor frequency $\sim\!9.6$ kHz, the light polarisation angle being set to $\theta=40^{\circ}$, and the strength of applied noise set to 140 mVrms such that $f_{\mathrm{noise}}=0.26$ Hz.